Develop the activity sequence model or models and determine the normal time for the following work activity: An assembly worker obtains six bolts in one hand from a bin located one step away on his worktable and puts one bolt each into six holes in the bracket within easy reach in front of him.
Answer #1

Activity sequence model = A3B0G3A0B0P6F0A0B0P0A0

Tn = 10 (3 + 3 + 6) = 10(12) = 120 TMU (4.3 sec)

Here,

1 TMU = 0.036 seconds

Tn = Normal Time